OpenStudy (anonymous):

You buy a house for $250,000 and it appreciates at a rate of 3% a year. How much will the house be worth after 3 years? Round to the nearest dollar.

OpenStudy (anonymous):

250000(1.03)^3 = 273182

OpenStudy (anonymous):

thanks! I have like 2more questions like this.. you think you can help?

OpenStudy (anonymous):

I can tyr!

OpenStudy (anonymous):

1) Find the amount of time it will take to double an investment of $400 with a rate of 4.5% compounded continuously. Round your answer to the nearest tenth. 2) Find the rate it would take to double an investment of $600 in 10 years if the rate is compounded continuously. Round your answer to the nearest tenth.

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

First, the formula for compounding continuously is: \[A = P \times e ^{0.045t}\]

OpenStudy (anonymous):

\[800 = 400 \times e ^{0.045t}\] \[2 = e ^{0.045t}\]\[\ln 2 = \ln e ^{0.045t}\]\[\frac{ \ln 2 }{ 0.045 } = t\]

OpenStudy (anonymous):

t = 15.4

OpenStudy (anonymous):

For the second one,\[1200 = 600 \times e ^{(r)(10)}\] \[2 = e ^{10r}\]\[\ln 2 = \ln e ^{10r}\]\[\frac{ \ln 2 }{ 10 } = r = 0.0693\]

OpenStudy (anonymous):

Thank you soo soo much!

OpenStudy (anonymous):

uw!